OpenStudy (anonymous):

Solve sin(7x)cos(x) + cos(7x)sin(x) = sin(4x)

OpenStudy (anonymous):

From 0 to 2pi

OpenStudy (anonymous):

LOL nice profile picture and name

OpenStudy (anonymous):

Haha thanks

OpenStudy (anonymous):

try sum and difference identities?

OpenStudy (anonymous):

Maybe, but I'm not supposed to know that for this course. Any other ideas?

OpenStudy (anonymous):

what course is that?

OpenStudy (anonymous):

Extension 1 Maths - Nsw Aus

OpenStudy (anonymous):

subtract sin 4x from both sides and graph it

OpenStudy (anonymous):

Graph it? Nah that would be cheating

OpenStudy (anonymous):

graph it and sum/difference identities are the only ways I can think of to solve it

OpenStudy (anonymous):

I'll give sum/difference a try

zepdrix (zepdrix):

Yah on the left side we can see the `Angle Sum Identity for Sine` expanded out. So if we work it backwards, we might be able to get where we need to go. We currently have something like this: \(\large \sin a\cos b+\sin b\cos a\) which will simplify down to \(\large \sin(a+b)\)

OpenStudy (anonymous):

Of course! Thanks

OpenStudy (anonymous):

Yeah that's done it

OpenStudy (anonymous):

yep, that's a sum identity as I said before. Thank you zepdrix